A motorist, traveling at the 120km/hr sees a broken down truck at the middle of the road and immediately applies his brake and comes to stop with uniform retardation in 20 secs. What distance does the car travel after the brakes were applied?


1
Expert's answer
2022-01-17T09:59:24-0500

Given:

"v_i=\rm 120 \:km/h=33.3\: m/s"

"v_f=0"

"t=20\:\rm s"


The distance traveled by a car

"d=\frac{v_i+v_f}{2}t=\frac{33.3+0}{2}*20=333\:\rm m"